OpenStudy (nuccioreggie):

Meg conducted a survey of four school cafeterias to find the number of students who like sandwiches for lunch. The results of her survey are recorded in the table below: School Cafeteria Survey School Total Number of Students in the Cafeteria Number of Students Who Like Sandwiches A 38 12 B 48 9 C 26 10 D 24 8 Which school has the greatest percentage of students who like sandwiches for lunch? School A School B School C School D

OpenStudy (nuccioreggie):

@pooja195

pooja195 (pooja195):

Easiest way to solve this is to turn the number of liked sandwiches into fractions

OpenStudy (nuccioreggie):

i thought it would be school C with 38%

pooja195 (pooja195):

\[\frac{ 12 }{ 38 }~\frac{ 9 }{ 48 }~\frac{ 10 }{ 26 }\frac{ 8 }{ 24 }\]

OpenStudy (nuccioreggie):

is C correct?

pooja195 (pooja195):

yes :)

OpenStudy (nuccioreggie):

thankyou (: